# Existence for Stable Rotating Star-Planet Systems

### v1

@@ 2.1 Variational formulation
    $W^{\infty}(\rho,\kappa)$ does not exceed the diameter of $\operatorname{spt}(\rho-\kappa)$
@@
    if $W^{\infty}(\rho,\kappa)<\delta$, each connected component of the $\delta$-neighbourhood of spt $\rho$ has the same mass for $\kappa$ as for $\rho$
@@
The advantage of the topology induced by Wasserstein $L^{\infty}$ metric is that a neighborhood of $\rho$ does not include such “portion-shifted-to-far-away” configurations mention above due to Lemma 2.9 (ii), while thanks to Lemma 2.9 (i), the neighborhood still contains all perturbations of the form $\rho+t\sigma$ for all $t$, where $\sigma$ is supported on a set which diameter is small enough, with $\int\sigma=0$. Hence this setting still allows one to legitimately consider variational derivatives (18) and consequently Euler-Lagrange equations (EL) (see also Remark 3.4.).
@@
According to Lemma 2.9, to prove $\rho(m)$ is a $W^{\infty}$ local minimizer (and hence solution to (EP’) due to Theorem 3.1), it suffices to show $\rho(m)$ is a constraint minimizer on $W_{m}$ such that its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. In fact, once we know $\rho(m)$’s support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$, one can also show $\rho(m)$ is a solution to (EP’), as well as most results in Theorem 3.1 (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) without specifying $\rho(m)$ is a $W^{\infty}$ local minimizer, see [15, Remark 3.10].
It motivates our choice to define $\Omega_{m}$ and $\Omega_{1-m}$ with their radii that increase as $\eta$ increases, and then show the size of stars or planets will not expand too much as $\eta$ increases. We will show those results for small $m$ later, inspired by McCann’s arguments for large $J$ [46, Section 6].
@@ 2.2 Statement of Main Results
    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int{\frac{\rho(m)(y)}{\left|{y-x}\right|}d^{3}y}}$. Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
Actually, the results can be true in a larger range of index $\gamma$ if we don’t require the support of the planet to shrink as the mass goes to 0.
@@
    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int{\frac{\rho(m)(y)}{\left|{y-x}\right|}d^{3}y}}$. Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
We provide the proof of Theorem 2.12 and Theorem 2.13 from Section 4 to Section 6. The general outline is as follows: Given Theorem 3.1, the remaining task is to demonstrate the existence of a Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizer. To achieve this, we first prove the existence of constrained minimizers $\rho(m)$ of $E_{J}(\rho)$ on $W_{m}$ when the mass ratio is sufficiently small in Section 4. In Section 5, we prove convergence of scaling densities in $L^{p}$, uniform $L^{\infty}$ bound, and then derives quantitative bounds on the supports of minimizers, . In Section 6 we estimate the center of mass separation and show that $\rho(m)$ can be situated within the interior of the considered domains we consider, that is, $\text{dist }\left(\text{spt}~\rho(m),~\mathbb{R}^{3}\backslash\left(\Omega_{m}\cup\Omega_{1-m}\right)\right)>0$, and then we verify $\rho(m)$ satisfies results in Theorem 2.12 and Theorem 2.13.
@@ 3 Preliminary Results
Recall that in the calculus of variations, if an energy functional admits a minimizer, then this minimizer satisfies Euler–Lagrange equation. When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, Theorem A.1 or Theorem 3.1. In this section, we review some variational‑based results and their extensions. In subsection 3.1, we discuss McCann’s work on Wasserstein $L^{\infty}$ local minimizers and existence of binary star system results. In subsection 3.2, we review some results for single-star system.
@@ 3.1 Results for 2-body systems
In this subsection, we review McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution with slight modification. We also make some remarks on them.
The following are McCann’s results on $W^{\infty}$ local minimizers.
Theorem 3.1 (Properties of $W^{\infty}$-Local Energy Minimizers [46, Section 2][15, Section 2]).
@@
    $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ if and only if $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and $v(x)=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\rho(m))}$, which implies $E(\rho,v)=E_{J}(\rho)$. It describes uniformly rotating fluid, in the sense that after defining $({\widetilde{\rho}}(t,x),v(t,x))=\left({\rho}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$. Notice that $E(\rho,v)=E(\widetilde{\rho},v)=E_{J}(\widetilde{\rho})=E_{J}(\rho)$.
@@
    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component;
@@
In Theorem 3.1 we only mention the case where $P(\rho)$ satisfies polytropic law (1), but the results hold true for a more general form of $P(\rho)$, see for instance [15, Theorem 2.25].
@@
Notice in Theorem 3.1, since the local minimizer $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\bar{x}(\rho)=(0,0,0)^{T}$ in (EL). However, one can generalize the case where the center of mass is not at the origin and $\rho$ is constrained minimizer instead of $W^{\infty}$ local minimizer, hence we retain the term $\bar{x}(\rho)$ in (EL) in Theorem 3.1 (iv) and (EL) holds on $\Omega_{i}$ instead of $\delta$-neighborhood of $\Omega_{i}$. This is consistent with the form in McCann’s paper [46, Section 6].
@@
McCann initially introduced and proved Theorem 3.1 in [46]. We later complemented this work in [15] by showing why (EP) also holds on the boundary of support of minimizer. Additionally, in [15] we verified the local minimizer $\rho$ has finite internal energy $U(\rho)<\infty$, hence $E_{J}$ is indeed differentiable at $\rho$ due to Lemma 2.6. Further discussion—including the reason why the local minimum energy is attained under uniform rotation (Theorem 3.1 (ii)), the interpretation of the stability result (Theorem 3.1 (viii)), the generalization of the fixed angular momentum from $J_{z}$ to $\mathbf{J}$, and the equivalence between (EP’) and (EL)—can be found in [46, 15, 31].
@@
Theorem 3.5 (Existence of Binary Stars [46, Theorem 6.1, Corollary 6.2]).
@@ 3.2 Results for Non-rotating Bodies
In this subsection, we introduce the existence and uniqueness of non-rotating solutions of Euler-Poisson equations, corresponding the single-star case. We also indicate some useful properties about the solutions.
@@
    $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave. In particular, for $m>0$, $e_{0}(m)<0$, $e_{0}^{\prime}\left(m^{+}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{+}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ and $e_{0}^{\prime}\left(m^{-}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{-}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ exists, and $e_{0}^{\prime}(m^{+})<e_{0}^{\prime}(m^{-})<0$;
@@
    $\left\|\sigma_{m}\right\|_{L^{\infty}}\leq C(m)$;
@@
Since $\sigma$ has mass 1 and satisfies Theorem 3.6 (iii), $\sigma\in L^{1}\cap L^{\infty}$, then $U(\sigma)$ is finite by interpolation inequality Proposition A.2. Hence the equation for $\lambda$ in Theorem 3.6 (vii) makes sense.
Remark 3.9.
In Theorem 3.6, we only mention the case where $P(\rho)$ satisfies polytropic law (1), but some results hold true for a more general form of $P(\rho)$, see for instance [4, 40, 46, 16]. Auchmuty and Beals initially concluded that (EP’) is satisfied in the region where $\rho>0$ [4, Theorem A]. We later improved the result in [16, Theorem 2.7] by showing the equation holds over the entire $\mathbb{R}^{3}$ space. Theorem 3.6 (ii) was stated without proof by McCann [46, Section 3], who mentioned it could be demonstrated using a method analogous to that in [40]—which is in the quantum mechanical framework. A proof for the polytropic case was later given in [16].
Note that without polytropic law assumption, the uniqueness of minimizer, the existence of $e^{\prime}_{0}(m)$ and the equality relations between $\lambda_{m}$, $U(\sigma$) and $e^{\prime}_{0}(m)$ in (vii) and (viii) are not guaranteed. Instead, we have $e_{0}^{\prime}\left(m^{+}\right)\leq\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)$. Note $e_{0}^{\prime}\left(m^{-}\right)<0$ if $m>0$ due to Theorem 3.6 (ii). Although uniqueness result in general case is unknown, uniqueness can still be established under assumptions slightly weaker than the polytropic law; see [46, 16].
Remark 3.10.
Intuitively, when the distance between two celestial bodies is sufficiently large, the kinetic energy and the graviational interaction energy becomes negligible. Consequently, $E_{J}$ will approach $e_{0}$ without equaling it. This intuition is indeed correct [16, Remark 2.43 and Remark 2.44]. In Section 5, we will also compare $e_{0}$ and the energy $E_{0}(\widetilde{\rho}_{m})$ for scaling density $\widetilde{\rho_{m}}$. This comparison allows us to apply concentration-compactness lemma in [42] and will also provide a clearer mathematical understanding of the relation between $e_{0}$ and $E_{J}$.
@@
Theorem 3.11 (Relations between Minimizers and Minimal Energies with Different Mass [16, Theorem 3.2]).
@@
Recall (19) that $E_{0}(\sigma)=U(\sigma)-\frac{G(\sigma,\sigma)}{2}=\int A(\sigma)-\frac{1}{2}\iint\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y$, where $A(\sigma)=\frac{K}{\gamma-1}\sigma^{\gamma}$. Then let $\sigma_{\widetilde{A},B}=\widetilde{A}\sigma(Bx)$, we can compute
$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int A\left(\sigma_{\widetilde{A},B}\right)$ (27)
$=\int\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}$
$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}d^{3}(Bx)$
@@
$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}d^{3}xd^{3}y$ (28)
$=\widetilde{A}^{2}B^{-5}\int\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}d^{3}(Bx)d^{3}(By)$
@@
$\int\sigma_{\widetilde{A},B}$ $=\widetilde{A}B^{-3}\int\sigma(Bx)d^{3}(Bx)$ (29)
$=\widetilde{A}B^{-3}\int\sigma$
From the scaling relations (27) (28) (29), together with the uniqueness result Theorem 3.6 (i), if we take $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, then we conclude the proof. For a complete discussion, one can see [16, Theorem 3.2]. ∎
@@
Thanks to Theorem 3.6, we know $\|\sigma\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ and spt $\sigma$ is contained in a ball of radius $R_{1}$, therefore, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{m}=\frac{C_{1}}{A}$ and spt $\sigma_{m}$ is contained in a ball of radius $R_{m}=BR_{1}$. Since $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, if we further assume $\gamma>2$, we know $\lim\limits_{m\rightarrow 0}A=+\infty$ and $\lim\limits_{m\rightarrow 0}B=0$, thus $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ and the size of $\sigma_{m}$’s support will go to 0 when $m\rightarrow 0$, with rates $m^{\frac{2}{3\gamma-4}}$ and $m^{\frac{\gamma-2}{3\gamma-4}}$ respectively. When $\gamma<2$, we still have $\lim\limits_{m\rightarrow 0}A=+\infty$, but $\lim\limits_{m\rightarrow 0}B=\infty$, and the size of $\sigma_{m}$’s support can go to $\infty$ when $m\rightarrow 0$ (flatten out or dispread). It coincides with a result mentioned in Lieb and Yau’s paper [40, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$. In their paper quantum mechanics (fermions case) is discussed.
@@
Now we know the minimizer satisfies (EL). Actually the converse also holds true: all radial solutions of (EL) are in fact minimizers of $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$ for some mass $m$. Furthermore, the minimizer (or radial solution to (EL)) can be uniquely parameterized by its central density, with each central density corresponding to a specific mass, see [16, Proposition 2.48]. Thanks to Theorem 3.6, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is actually the central density of $\sigma_{m}$, thus Theorem 3.11 shows as the central density decreases, the mass also decreases. Moreover, it gives the decay rate. For more general pressure $P(\rho)$, this decreasing relationship is also established in [16, Corollary 2.41].
@@
Therefore, we can further discuss the relationship of variational derivatives (see Lemma 2.6) between densities with different mass (not necessary to be minimizers).
Proposition 3.14 (Relations of Variational Derivatives Between Densities with Different Mass [16, Proposition 3.4]).
Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies $E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}$ $E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$, where $\rho(x):=\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, $\rho$ has mass 1.

### v2

@@ 2.1 Variational formulation
    $W^{\infty}(\rho,\kappa)$ does not exceed the diameter of $\operatorname{spt}(\rho-\kappa)$;
@@
    if $W^{\infty}(\rho,\kappa)<\delta$, each connected component of the $\delta$-neighbourhood of spt $\rho$ has the same mass for $\kappa$ as for $\rho$. Here the $\delta$-neighbourhood of some $\Omega\subset\mathbb{R}^{3}$ is defined as $\bigcup_{y\in\Omega}\left\{x\in\mathbb{R}^{3}\mid|x-y|<\delta\right\}$.
@@
The advantage of the topology induced by Wasserstein $L^{\infty}$ metric is that a neighborhood of $\rho$ does not include such “portion-shifted-to-far-away” configurations mention above due to Lemma 2.9 (ii), while thanks to Lemma 2.9 (i), the neighborhood still contains all perturbations of the form $\rho+t\sigma$ for all $t$, where $\sigma$ is supported on a set whose diameter is small enough, with $\int_{\mathbb{R}^{3}}\sigma\,dx=0$. Hence this setting still allows one to legitimately consider variational derivatives (19) and consequently Euler-Lagrange equations (EL) and Euler-Poisson equations (EP’) (see also Remark 3.4).
@@
According to Lemma 2.9, to prove $\rho(m)$ is a $W^{\infty}$ local minimizer (and hence a solution to (EP’) due to Theorem 3.1), it suffices to show $\rho(m)$ is a constraint minimizer on $W_{m}$ such that its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. In fact, once we know $\rho(m)$’s support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$, one can also show $\rho(m)$ is a solution to (EP’), as well as most results in Theorem 3.1 (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) without specifying $\rho(m)$ is a $W^{\infty}$ local minimizer, see [10, Section 3].
It motivates our choice to define $\Omega_{m}$ and $\Omega_{1-m}$ with their radii that increase as $\eta$ increases, and then show the size of stars or planets will not expand too much as $\eta$ increases. We will show those results for small $m$ later, inspired by McCann’s arguments for large $J$ [24, Section 6].
@@ 2.2 Statement of Main Results
    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int_{\mathbb{R}^{3}}{\frac{\rho(m)(y)}{\left|{y-x}\right|}\,dy}}$. Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
Actually, the results can be true in a larger range of index $\gamma$ if we do not require the support of the planet to shrink as the mass goes to 0.
@@
    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int_{\mathbb{R}^{3}}{\frac{\rho(m)(y)}{\left|{y-x}\right|}\,dy}}$. Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
We provide the proof of Theorem 2.12 and Theorem 2.13 from Section 4 to Section 6. The general outline is as follows: Given Theorem 3.1, the remaining task is to demonstrate the existence of a Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizer. To achieve this, we first prove the existence of constrained minimizers $\rho(m)$ of $E_{J}(\rho)$ on $W_{m}$ when the mass ratio is sufficiently small in Section 4. In Section 5, we prove convergence of scaling densities in $L^{p}$, uniform $L^{\infty}$ bound, and then derive quantitative bounds on the supports of minimizers. In Section 6 we estimate the center of mass separation and show that $\rho(m)$ can be situated within the interior of the considered domains we consider, that is, $\text{dist }\left(\text{spt}~\rho(m),~\mathbb{R}^{3}\backslash\left(\Omega_{m}\cup\Omega_{1-m}\right)\right)>0$, and then we verify $\rho(m)$ satisfies results in Theorem 2.12 and Theorem 2.13.
@@ 3 Preliminary Results
Recall that in the calculus of variations, if an energy functional admits a minimizer, then this minimizer satisfies Euler–Lagrange equation. When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, Theorem 3.1 or [4, Section 2]. In this section, we review some variational‑based results and their extensions. In subsection 3.1, we discuss McCann’s work on Wasserstein $L^{\infty}$ local minimizers and existence of binary star system results. In subsection 3.2, we review some results for single-star systems.
@@ 3.1 Results for 2-body systems
In this subsection, we review McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution with slight modifications. We also make some remarks on them.
The following are McCann’s results for $W^{\infty}$ local minimizers.
Theorem 3.1 (Properties of $W^{\infty}$-Local Energy Minimizers [24, Section 2][10, Section 2]).
@@
    $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ if and only if $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and $v(x)=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\rho(m))}$, which implies $E(\rho,v)=E_{J}(\rho)$. It describes uniformly rotating fluid, in the sense that after defining $({\widetilde{\rho}}(t,x),v(t,x))=\left({\rho}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$. In particular, $E(\rho,v)=E(\widetilde{\rho},v)=E_{J}(\widetilde{\rho})=E_{J}(\rho)$.
@@
    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component. Here $\cdot$$_{+}$ is the nonnegative (positive) part function defined as $\lambda]_{+}:=\max\{\lambda,0\}$;
@@
In Theorem 3.1 we only mention the case where $P(\rho)$ satisfies polytropic law (1), but the results hold true for a more general form of $P(\rho)$, see for instance [10, Section 2].
@@
Notice in Theorem 3.1, since the local minimizer $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\bar{x}(\rho)=(0,0,0)^{T}$ in (EL). However, one can generalize to the case where the center of mass is not at the origin, hence we retain the term $\bar{x}(\rho)$ in (EL) in Theorem 3.1 (iv). This is consistent with the form in McCann’s paper [24, Section 6]. Furthermore, if $\rho$ is a constrained minimizer instead of a $W^{\infty}$ local minimizer, (EL) still holds on $\Omega_{i}$.
@@
McCann initially introduced and proved Theorem 3.1 in [24]. We later complemented this work in [10] by showing why (EP) also holds on the boundary of the support of minimizer. Additionally, in [10] we verified the local minimizer $\rho$ has finite internal energy $U(\rho)<\infty$, hence $E_{J}$ is indeed differentiable at $\rho$ due to Lemma 2.6. Further discussion—including the reason why the local minimum energy is attained under uniform rotation (Theorem 3.1 (ii)), the interpretation of the stability result (Theorem 3.1 (viii)), the generalization of the fixed angular momentum from $J_{z}$ to $\mathbf{J}$, and the equivalence between (EP’) and (EL)—can be found in [24, 10, 17].
@@
Theorem 3.5 (Existence of Binary Stars [24, Theorem 6.1, Corollary 6.2]).
@@ 3.2 Results for Non-rotating Bodies
In this subsection, we introduce the existence and uniqueness of non-rotating solutions of Euler-Poisson equations, corresponding to the single-star case. We also indicate some useful properties about the solutions.
@@
    $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave. In particular, for $m>0$, $e_{0}(m)<0$, $e_{0}^{\prime}\left(m^{+}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{+}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ and $e_{0}^{\prime}\left(m^{-}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{-}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ exists, and we have $e_{0}^{\prime}(m^{+})\leq e_{0}^{\prime}(m^{-})<0$;
@@
    $\left\|\sigma_{m}\right\|_{L^{\infty}}\leq C(m)$, and $U(\sigma_{m})<\infty$ due to Remark 2.5;
@@
In Theorem 3.6, we only mention the case where $P(\rho)$ satisfies polytropic law (1), but some results hold true for a more general form of $P(\rho)$, see for instance [4, 20, 24, 11]. Auchmuty and Beals initially concluded that (EP’) is satisfied in the region where $\rho>0$ [4, Theorem A]. We later improved the result in [11, Section 2] by showing the equation holds over the entire $\mathbb{R}^{3}$ space. Theorem 3.6 (ii) was stated without proof by McCann [24, Section 3], who mentioned it could be demonstrated using a method analogous to that in [20]—which is in the quantum mechanical framework. A proof for the polytropic case was later given in [11].
Note that without polytropic law assumption, the uniqueness of minimizer, the existence of $e^{\prime}_{0}(m)$, and the equality relations between $\lambda_{m}$, $U(\sigma$) and $e^{\prime}_{0}(m)$ in (vii) and (viii) are not guaranteed. Instead, we have $e_{0}^{\prime}\left(m^{+}\right)\leq\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)$. Note $e_{0}^{\prime}\left(m^{-}\right)<0$ if $m>0$ due to Theorem 3.6 (ii). Although uniqueness result in the general case is unknown, uniqueness can still be established under assumptions slightly weaker than the polytropic law; see [24, 11].
@@
Theorem 3.9 (Relations between Minimizers and Minimal Energies with Different Mass [11, Section 3]).
@@
Recall (20) that $E_{0}(\sigma)=U(\sigma)-\frac{G(\sigma,\sigma)}{2}=\int_{\mathbb{R}^{3}}A(\sigma)\,dx-\frac{1}{2}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy$, where $A(\sigma)=\frac{K}{\gamma-1}\sigma^{\gamma}$. Then let $\sigma_{\widetilde{A},B}=\widetilde{A}\sigma(Bx)$, we can compute
$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int_{\mathbb{R}^{3}}A\left(\sigma_{\widetilde{A},B}\right)\,dx$ (28)
$=\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}\,dx$
$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}\,d(Bx)$
@@
$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}\,dx\,dy$ (29)
$=\widetilde{A}^{2}B^{-5}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}\,d(Bx)\,d(By)$
@@
$\int_{\mathbb{R}^{3}}\sigma_{\widetilde{A},B}\,dx$ $=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma(Bx)\,d(Bx)$ (30)
$=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma\,dx$
From the scaling relations (28) (29) (30), together with the uniqueness result Theorem 3.6 (i), if we take $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, then we conclude the proof. For a complete discussion, one can see [11, Section 3]. ∎
@@
Thanks to Theorem 3.6, we know $\|\sigma\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ and spt $\sigma$ is contained in a ball of radius $R_{1}$, therefore, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{m}=\frac{C_{1}}{A}$ and spt $\sigma_{m}$ is contained in a ball of radius $R_{m}=BR_{1}$. Since $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, if we further assume $\gamma>2$, we know $\lim\limits_{m\rightarrow 0}A=+\infty$ and $\lim\limits_{m\rightarrow 0}B=0$, thus $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ and the size of $\sigma_{m}$’s support will go to 0 when $m\rightarrow 0$, with rates $m^{\frac{2}{3\gamma-4}}$ and $m^{\frac{\gamma-2}{3\gamma-4}}$ respectively. When $\gamma<2$, we still have $\lim\limits_{m\rightarrow 0}A=+\infty$, but $\lim\limits_{m\rightarrow 0}B=\infty$, and the size of $\sigma_{m}$’s support can go to $\infty$ when $m\rightarrow 0$ (flatten out or dispread). It coincides with a result mentioned in Lieb and Yau’s paper [20, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$. In their paper quantum mechanics (fermions case) is discussed.
@@
Now we know the minimizer satisfies (EL). Actually the converse also holds true: all radial solutions of (EL) are in fact minimizers of $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$ for some mass $m$. Furthermore, the minimizer (or radial solution to (EL)) can be uniquely parameterized by its central density, with each central density corresponding to a specific mass, see [11, Section 2]. Thanks to Theorem 3.6, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is actually the central density of $\sigma_{m}$, thus Theorem 3.9 shows as the central density decreases, the mass also decreases. Moreover, it gives the decay rate. For more general pressure $P(\rho)$, this decreasing relationship is also established in [11, Section 2].
@@
Therefore, we can further discuss the relationship between variational derivatives (see Lemma 2.6) of densities with different mass (not necessarily to be minimizers).
Proposition 3.12 (Relations Between Variational Derivatives of Densities with Different Mass [11, Section 3]).
Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies
$E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
where $\rho(x):=\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$. In particular, $\rho$ has mass 1.